The minimum angular magnification of an astronomical telescope is 10 the length of it's tube is 44cm determine the focal length of the lenses in normal adjustment

Grade:12

2 Answers

Aditya Gupta
4 years ago
kindly follow the link:
if u still have any doubt feel free 2 ask :))
Vikas TU
4 years ago
Dear student 
The ans given by ADITYA  is incorrect , It is not as per question.
Magnifying power of astronomical telescope = Focal length of Objective/Focal length of eye piece M = f0/fe ....(i) Distance between two lenses = Focal length of eye - piece + Focal length of the objective.
d = fe + f0 
M = 10 
d = 44 
f0/fe = 10 
f0 = 10fe 
11fe = 44 
fe = 4 cm 
f0 = 44 cm 
fe = 4 cm 
Hope this will help 
Good Luck